# Sensor-Based Distributionally Robust Control for Safe Robot Navigation in Dynamic Environments

### v1

@@ 9.2 Simulated Dynamic Environments
Figure 10: Evaluation of distance to obstacles and robot velocity profile over time in a simulated dynamic environment in Fig. 5(b). The timestamps corresponding to the snapshots in Fig. 9 are indicated with vertical dotted lines, demonstrating the robot’s adaptive responses to dynamic obstacles.
In this section, we evaluate our CLF-DR-CBF QP approach in safe navigation in dynamic environments, cf. Fig. 5(b). Using a social force model (Helbing and Molnar, 1995; Moussaïd et al., 2010), we simulate a scenario within Gazebo that mirrors real-world environments with pedestrians. This simulation aims to reflect the complexities that a mobile robot faces in human-populated areas. Due to limitations of the Gazebo simulator, the position and velocity information of the moving agents are assumed known and provided by Gazebo. In our experiments in Sec. 9.3, these dynamic obstacles are estimated by the onboard sensing. The static elements of the environment are unknown and the robot relies on its LiDAR measurements with noise $\sigma=0.01$ to ensure safety.
In the dynamic simulation shown in Fig. 9, the robot starts at $(0,0)$ and is tasked to sequentially visit two waypoints before reaching a designated goal at $(3,0)$. This task is complicated by the presence of three pedestrians, requiring the robot to dynamically adjust its trajectory to avoid collisions while still aiming to complete its mission efficiently. Note that the $A^{*}$ algorithm operates independently of the pedestrian motion, and the real-time pedestrian avoidance relies on our CLF-DR-CBF QP formulation. This separation emphasizes the role of our approach in ensuring safety in dynamic environments, where pedestrian movement and noisy LiDAR measurements are challenging to consider in the high-level path planning phase.
In Fig. 9(a), at $t=3.4$s, the robot encounters a pedestrian on a collision course with its planned path to the first waypoint at the top right. With our CLF-DR-CBF QP controller, the robot employs a defensive maneuver. This adjustment shows the methodology’s capability to anticipate potential hazards and react accordingly.
@@
Furthermore, Fig. 10 provides the robot’s distance to obstacles and its velocity profiles over time. The plot confirms that the robot maintains safe distance from the obstacles, including pedestrians, while exhibiting smooth linear and angular velocities.
@@ 9.3 Real-World Experiments
We demonstrate the performance of our CLF-DR-CBF QP approach in a cluttered lab environment, where the robot relies solely on LiDAR measurements to navigate around various challenges, such as thin chair legs, narrow passages with pedestrians, and approaching pedestrians. In contrast to the CLF-GP-CBF SOCP formulation, which relies on GP regression to construct CBFs and cannot handle dynamic environments effectively, our CLF-DR-CBF QP formulation maintains safety while solely depending on noisy LiDAR measurements. Fig. 11 shows these challenging scenarios. The bottom plot in Fig. 11 presents the distance to the obstacles and the robot’s velocity profile over time, highlighting the robot’s ability to maintain a safe distance while efficiently navigating towards its goal. For more details, please refer to the accompanying videos on the project webpage [^2] https://existentialrobotics.org/DR_Safe_Navigation_Webpage/.
@@
10 Conclusion
We introduced a novel strategy for ensuring safety of mobile robots navigating autonomously in unknown dynamically changing environments. Our distributionally robust control barrier function formulation leverages sensor measurements directly, eliminating the need for precise CBF estimation, which may be slow and inaccurate in dynamic environments. By combining the DR-CBF with a control Lyapunov function design for path tracking, we developed a CLF-DR-CBF quadratic program that enables safe autonomous navigation for robots with control-affine dynamics. Our approach underscores the efficiency and effectiveness of employing fast sensor-based DR-CBF constraints to handle measurement uncertainty and the complexities of real-world environments. Our experiments demonstrate the ability of our CLF-DR-CBF control method to respond to fast environment changes, while maintaining progress towards a desired navigation goal.
Our results suggest that further exploration into adaptive and sensor-based techniques for robot control may lead to significant progress in deploying reliable autonomous robot systems in the real world. Future work will focus on extending this methodology to more complex robot systems, such as robot arms and humanoids, further bridging the gap between theoretical safety assurances and practical considerations.

### v2

@@ 9.2 Simulated Dynamic Environments
Table 4: Performance metrics for dynamic environments over 1000 trials. The metrics include success rate, stuck rate, collision rate, and task completion time (mean ± std). The CLF-DR-CBF QP outperforms both baselines in terms of success rate and collision avoidance, demonstrating its robustness in dynamic settings.
Method Success Rate (%) Stuck Rate (%) Collision Rate (%) Time (s)
CLF-DR-CBF QP 93.2 5.1 1.7 $10.7\pm 2.2$
CLF-GP-CBF SOCP 60.5 36.3 3.2 $13.6\pm 2.9$
Nominal CLF-CBF QP 61.7 8.5 29.8 $10.1\pm 2.1$
@@
The dynamic environment simulations are conducted in Gazebo (cf. Fig. 5(b)), designed to mimic real-world scenarios with pedestrians modeled using the social force model (Helbing and Molnar, 1995; Moussaïd et al., 2010).
Hypothesis: Our CLF-DR-CBF QP approach will outperform baseline methods in handling time-varying constraints under noisy conditions. Specifically, we expect our method to achieve higher success rates, lower collision rates, and efficient task completion times, due to its ability to incorporate sensor noise and localization uncertainty directly into the control formulation.
Setup: The robot starts at $(0,0)$ with an initial orientation of $0^{\circ}$, and the goal locations are randomly placed at least $6$ meters away. Pedestrians are also randomly positioned in the environment, with velocities bounded by $B=1$ m/s. Both the static and dynamic elements of the environment are unknown, and the robot relies on noisy LiDAR measurements (Gaussian noise with $\sigma=0.05$) for collision avoidance. In all simulations below, the $A^{*}$ planning algorithm operates independently of the pedestrian motion, and the real-time pedestrian avoidance relies on our CLF-DR-CBF QP formulation (or the two baseline approaches). We conducted 1000 trials for each method, measuring metrics such as success rate, stuck rate, collision rate, and average task completion time. Success is defined as reaching the goal while maintaining safety, while a trial is considered stuck if the robot fails to reach the goal within 20 seconds.
Results and discussion: Table 4 summarizes the quantitative results. The proposed CLF-DR-CBF QP approach achieves the highest success rate (93.2%) and the lowest collision rate (1.7%) among the three methods. By directly handling sensor noise through its distributionally robust formulation, the CLF-DR-CBF QP method ensures safety while maintaining efficient task completion times. The GP-CBF method exhibits a lower success rate and higher stuck rate due to its reliance on the GP-SDF map, which becomes computationally expensive and less reliable in dynamic environments. The Nominal CLF-CBF QP approach suffers from the highest collision rate (29.8%), highlighting its limitations in handling dynamic obstacles with sensor noise.
We next present some qualitative results in Fig. 8 in the same dynamic environment. The robot is tasked to sequentially visit two waypoints before reaching a designated goal at $(3,0)$. In Fig. 8(a), at $t=3.4$s, the robot encounters a pedestrian on a collision course with its planned path to the first waypoint at the top right. With our CLF-DR-CBF QP controller, the robot employs a defensive maneuver. This adjustment shows the methodology’s capability to anticipate potential hazards and react accordingly.
@@ 9.3 Real-World Experiments
(a) Robot Shape 1
(b) Robot Shape 2
Figure 9: Robot shapes used in real-world experiments.
Figure 10: Comparison of trajectories for the three robot shapes.
@@
Setup: The experiments took place in a cluttered lab environment designed to test various challenging scenarios. The robot relied solely on noisy LiDAR measurements for navigation. We first tested the robot in an area of the lab with static obstacles, using three different shapes: the original shape (Fig. 1(a)), Shape 1 (Fig. 9(a)), and Shape 2 (Fig. 9(b)). For quantitative evaluation, we ran 50 trials per shape in environments populated with randomly placed obstacles (e.g., cubes and pyramids) and three randomly walking pedestrians.
Results and discussion: Figure 10 shows the trajectories of the robot with the three shapes navigating the same environment by following a pre-planned path. Notably, the planned path was generated assuming the nominal robot shape and did not account for the differences introduced by Shape 1 and Shape 2. Consequently, the original shape demonstrates a minimal deviation from the planned path, whereas Shape 1 exhibits the largest deviations due to its wider, asymmetrical design.
Table 5 summarizes the results, showing the success, stuck, and collision rates for the three robot shapes under the proposed CLF-DR-CBF QP formulation. Our approach achieved high success rates across all robot shapes, demonstrating its robustness to variations in geometry and dynamic obstacles in the environment. In our evaluation, a robot was considered “stuck” if it did not reach its goal within 60 seconds.
When the robot shape becomes larger (Shapes 1 and 2), it is more likely to get stuck due to limited maneuverability in tight spaces. This is particularly evident for Shape 1, where the asymmetry makes the robot significantly wider on one side, further complicating its ability to bypass obstacles. Additionally, the larger size of Shapes 1 and 2 reduces their ability to navigate around obstacles within the required time, leading to reduced success rates. Despite these challenges, all three shapes achieved a 0% collision rate, demonstrating the effectiveness of our CLF-DR-CBF QP formulation in maintaining safe navigation.
Table 5: Performance metrics for real-world experiments with different robot shapes. Metrics include success rate, stuck rate, and collision rate over 50 trials per shape.
Shape Success(%) Stuck (%) Collision(%)
Original 100 0 0
Shape 1 84 16 0
Shape 2 90 10 0
Next, we demonstrated the performance of our CLF-DR-CBF QP formulation using the original robot shape in a full-lab navigation task. The robot successfully handled various real-world challenges, such as thin chair legs, narrow passages with pedestrians, and approaching pedestrians (Fig. 11). In contrast to the CLF-GP-CBF SOCP formulation, which relies on GP regression to construct CBFs and cannot handle dynamic environments effectively, our CLF-DR-CBF QP formulation maintains safety while solely depending on noisy LiDAR measurements. The bottom plot in Fig. 11 presents the distance to the obstacles and the robot’s velocity profile over time, highlighting the robot’s ability to maintain a safe distance while efficiently navigating towards its goal. For more details, please refer to the accompanying videos on the project webpage [^2] https://existentialrobotics.org/DRO_Safe_Navigation/.
@@
10 Conclusion and Future Work
We introduced a novel strategy for ensuring safety of mobile robots navigating autonomously in unknown dynamically changing environments. Our distributionally robust control barrier function formulation leverages sensor measurements and state estimates directly, eliminating the need for precise knowledge of CBFs, which may be slow and inaccurate to obtain in dynamic environments. By combining the DR-CBF with a control Lyapunov function for path tracking, we developed a CLF-DR-CBF quadratic program that enables safe autonomous navigation for robots with control-affine dynamics and arbitrary shape. Our approach underscores the efficiency and effectiveness of employing sensor-based DR-CBF constraints to handle uncertainty in measurements and state estimates. The simulation and experiment results suggest that further exploration into integrating sensor measurements and perception estimates directly into the robot planning and control methods may lead to significant progress in deploying reliable autonomous robot systems in the real world.
While our methodology demonstrates robustness and effectiveness, it has certain limitations. Sections 6 present a general framework for control-affine systems, relying on the assumption that valid CBFs with relative degree 1 exist for these systems. However, identifying such functions can be nontrivial, particularly for complex or high-dimensional systems. To provide a practical demonstration, we employed a signed distance function as a CBF for a differential-drive robot in Section 8 and conducted our evaluations on this specific model.
Future work will aim to extend this methodology to more complex robot systems, such as table-top and mobile manipulators. Additionally, we plan to explore high-order control barrier functions to address challenges associated with designing valid CBFs, further bridging the gap between theoretical safety assurances and practical deployments in real-world scenarios.